OpenStudy (legomyego180):

A 200 liter tank initially full of water develops a leak at the bottom. Given that 20% of the water leaks out in the first 10 minutes, find the amount of water left in the tank t minutes after the leak develops if the water drains off at a rate that is proportional to the amount of water present.

OpenStudy (legomyego180):

@Laynalove

OpenStudy (legomyego180):

@zzr0ck3r

OpenStudy (legomyego180):

@jim_thompson5910

OpenStudy (legomyego180):

@thephysicsman

OpenStudy (thephysicsman):



OpenStudy (thephysicsman):

I think we have to use related rates with this problem

OpenStudy (legomyego180):

nah, I wish. This is for my chapter on differential equations. I need to use the formula for exponential growth and decay

OpenStudy (thephysicsman):

woah, lol what class is that for?

OpenStudy (legomyego180):

just Cal II

OpenStudy (legomyego180):

its not actually differential equations lol

OpenStudy (thephysicsman):

You said something about exponential growth and decay \[\frac{ dA }{ dt } = e^{-kt}*A_{0}\]

OpenStudy (legomyego180):

\[P=A _{0}e ^{-kt}\] Yea thats the one

OpenStudy (thephysicsman):

@legomyego180 what else were you thinking?

OpenStudy (legomyego180):

I have a hunch on the equation? I think I need to solve for k, the decay constant first.

OpenStudy (legomyego180):

\[160=200e^{-k10}\]

OpenStudy (legomyego180):

how does that look to you?

OpenStudy (legomyego180):

Im getting \[k=-\frac{ \ln \frac{ 4 }{ 5 } }{ 10 }\] Which we can use to plug into our original equation to solve for t?

OpenStudy (thephysicsman):

\[\frac{ 4 }{ 5 } = e^{-k*10}\] \[\ln(\frac{ 4 }{ 5 }) = -10k \] \[\frac{ \ln(\frac{ 4 }{ 5 }) }{ -10 } = k \]

OpenStudy (thephysicsman):

yeah I agree :) but I feel there should be more to this problem?

OpenStudy (legomyego180):

Yea we need to plug k back into the original equation to solve for t but Im not sure what I need to do with the inital values variable and end value variable

OpenStudy (legomyego180):

@zepdrix @UnkleRhaukus you guys around?

OpenStudy (legomyego180):

@mww

OpenStudy (sdfgsdfgs):

this is VERY similar to a Q u posted a few days ago: http://openstudy.com/study#/updates/577474a5e4b0140fe65398af u could easily re-use the same equation and solve for the different constants in this case.

OpenStudy (mww):

they just need you to put back the k into the original equation (find the amount at t minutes)

zepdrix (zepdrix):

The rate of change is proportional to the amount at t time, \(\large\rm A'=kA\) and solving for A gives us the equation for growth/decay that we're used to seeing, \(\large\rm A=A_o e^{kt}\) Ya looks like you guys got everything figured out :) \(\large\rm k=\dfrac{\ln.8}{10}\) Plugging it in gives you,\[\large\rm A(t)=200e^{\left(\frac{\ln.8}{10}~t\right)}\]

zepdrix (zepdrix):

That's all they wanted.

zepdrix (zepdrix):

You don't need a negative on the k when you setup the equation. The negative will come out naturally. But I guess it's ok if you do.

OpenStudy (mww):

your initial amount is already included as P(0)= A = 200 when t = 0.

OpenStudy (legomyego180):

Gotcha, I think I actually have a fighting chance on this subject on my test tomorrow

OpenStudy (mww):

In fact if they say that a proportion of the original was left at time t,, then you can always write: \[\frac{ P(t) }{ A } = C = e^{-kt}\] immediately on LHS for some constant C being the proportion.